If the total mass hanging on one end of the string that creates the tension in the string is 1.20 kg, what is the mass density (mass per unit length) ? of the string? L = 0.800 m, frequency of 24.0 Hz.
Answer #1

mass =1.2kg

l = 0.8m

mass density = m/l

= 1.2/0.8

= 1.5 kg/m
Answer #2

The gravity force on the mass supplies the tension

m . g = 1.2 * 9.81 = 11.772 N

m = 1.2 kg

Increasing the mass will increase the tension, which will increase the velocity. With a fixed frequency this means that the wavelength must increase.
The maximum wavelength possible is 1.6 m

so the maximum speed ( to produce a coherent standing wave ) = 24 . 1.6 = 38.4 m/s

38.4 = ? T / mu

mu^2 = 11.772/1474.56

mu = 0.089 kg/m
Answer #3

frequency of mass loaded string is given by

f = \frac{1}{2L}\sqrt{\frac{T}{\mu }}

here tension T= Mg = 1.2 * 9.8 =11.76N and length of string L = 0.800m, frequency f = 24Hz

so 24 = \frac{1}{2*0.800}\sqrt{\frac{11.76}{\mu }}

so linear density \mu = 0.007975kg/m